Use the method of undetermined coefficients to solve the given nonhomogeneous system. X' = 4 1 3 9 6 X + −5 2 et
Answer #1

sol. X: (41)x+1)et for complementry solution, we use to find eigen values & eigen vecte of A= / 4 values 4 egen llet) (96) G→ 3) (3) = (; 3)-(:)+(-det = 45+49-5 C = 94 +6C2+2 gcg + 50 = -2 36 - Required sol is